OpenStudy (thats_kyy):

A system of equations is shown below: n = 3m + 6 n − 2m = 2 What is the solution, in the form (m, n), to the system of equations? (−4, −6) (−5, −9) (2, 6) (1, 8)

OpenStudy (thats_kyy):

@TheSmartOne

OpenStudy (thats_kyy):

Help Please

jimthompson5910 (jim_thompson5910):

n - 2m = 2 3m+6 - 2m = 2 ... replace n with 3m+6 (valid because n = 3m+6 is the first equation) 3m - 2m + 6 = 2 m+6 = 2 m = ???

OpenStudy (thats_kyy):

m = -4

OpenStudy (thats_kyy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes, correct

jimthompson5910 (jim_thompson5910):

plug that into n = 3m+6 to find the value of n

OpenStudy (thats_kyy):

Oh Thankss! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem